10cm cube of CO is mixed with 100cm cube of air contain 21% oxygen. If all the volume are measured at S.T. P calculate the volume of the resulting gases (air=100)
 

Grade:11

1 Answers

Vikas TU
7 years ago
moles of CO at STP = (10/1000)/22.4 => 0.01/22.4 = 0.000446
moles of O2 at STP = (21/1000)/22.4 = > 0.021/22.4 = 0.0009375
Total moles = 0.0013835
Volume at STP = > 0.0013835*22.4 => 0.0309904 litre or 30.9904 cm^3